# Infinitely many non-collapsed steady Ricci solitons on complex line bundles

### v1

@@ Abstract.
We construct a continuous 3-parameter family of non-shrinking Ricci solitons complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{2m+1}$, where the base space is not necessarily Kähler–Einstein. Each $\mathcal{O}(k)$ with $k\in[3,2m+1]$ admits at least one asymptotically conical (AC) Ricci-flat metric in this family. For each $\mathcal{O}(k)$ with $k\geq 3$, the family includes infinitely many asymptotically paraboloidal (AP) steady Ricci soliton.
2020 Mathematics Subject Classification: 53E20 (primary), 53C25 (secondary). Keywords: Ricci soliton, cohomogeneity one metric. The author is supported by the NSFC (No. 12071489, No. 12301078), the Natural Science Foundation of Jiangsu Province (BK-20220282), and XJTLU Research Development Fund (RDF-21-02-083).
@@ 1. Introduction
Ricci solitons are candidates for blow-up singularities of the Ricci flow. Specifically, a shrinking Ricci soliton is a candidate for a Type I singularity while a steady Ricci soliton is a candidate for a Type II singularity. From Perelman’s no-collapsing theorem [Per02], Type II singularity must be non-collapsed.
Complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{n}$ provide a rich source of examples for gradient Ricci solitons. The principal orbit of $\mathcal{O}(k)$ is the lens space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$, a circle bundle over $\mathbb{CP}^{n}$. The integer $k$ classifies the circle bundle, corresponding to a multiple of the indivisible element in $H^{2}(\mathbb{CP}^{n},\mathbb{Z})$. This setup has led to numerous examples of Ricci solitons with various asymptotic behaviors. The Bérard-Beregery metric on $\mathcal{O}(n+1)$ is an AC Ricci-flat metric; see [BB82]. By [Cao96], it is known that $\mathcal{O}(n+1)$ admits a 1-parameter family of steady Kähler–Ricci solitons. These solitons have collapsed volume growth. In particular, they are asymptotically cigar-paraboloidal (ACP). Expanding Kähler–Ricci solitons on $\mathcal{O}(k)$ with $k>n+1$ are found in [FIK03]. The $U(n+1)$-invariant ansatz is further generalized to complex line bundles over products of Fano manifolds in [DW11]. Recently, new steady Ricci solitons were found on $\mathcal{O}(k)$ independently by [Win23], [Sto24] and [App17]. It was proved in [App17] that each $\mathcal{O}(k)$ with $k>n+1$ admits a non-collapsed AP steady Ricci soliton.
@@
The group action descends to $\mathbb{CP}^{2m+1}$, where the isotropy subgroup of $[0:...:0:1]$ is
@@
Since the principal orbit is locally isometric to $\mathbb{S}^{4m+3}$, the cohomogeneity one dynamical system on $\mathcal{O}(k)$ are the same with various values of $k$. For $k\in\{1,2\}$, we construct a compact invariant set to obtain the following theorem.
@@
On each $\mathcal{O}(k)$ with $13\leq k\leq 2m+$ 1, there exists at least one AC Ricci-flat metric, whose base of the cone is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. On each $\mathcal{O}(k)$ with $k\geq 3$, there exist infinitely many AP steady Ricci solitons whose base of the paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$.
This article is structured as the following. We begin by presenting the cohomogeneity one Ricci soliton equation and introducing a new coordinate transformation, differing from that in [Chi24], to establish local existence. For global analysis, we first construct an invariant compact set to prove Theorem 1.1. Next, we generalize [App17, Theorem 5.1] and prove Theorem 1.2-1.3. Finally, we investigate the asymptotics for these Ricci solitons and prove Theorem 1.4.
@@ 2. Cohomogeneity one equation
The geometric setup is similar to the one in [Chi24]. Specifically, let $Q$ be the inner product on $\mathfrak{g}$ that generates the round metric on $\mathbb{S}^{4m+3}$. We have a $Q$-orthogonal decomposition of the isotropy representation $\mathfrak{g}/\mathfrak{k}$ into three irreducible summands. Consider the ansatz
@@
We set $f(0)=0$ without loss of generality. By [Buz11], the potential function $f$ is an even function in $t$ near 00. As the $\mathbb{S}^{1}$-fiber collapses at the singular orbit, we have $\mathrm{tr}(L)=\frac{1}{t}+O(t)$ as $t\to 0$. Therefore, we have $2\ddot{f}(0)=C$ by(2.3). As the radius for the $\mathbb{S}^{1}$-fiber of $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ is $\frac{1}{k}$, we require $\dot{a}(0)=k$ for smooth extension and we consider the initial condition
@@
(2.10) $\begin{split}\begin{bmatrix}Q\\ H\end{bmatrix}^{\prime}=\begin{bmatrix}2Q(G-\frac{\epsilon}{2}W)+\epsilon(H-1)W\\ (H-1)\left(G-\frac{\epsilon}{2}W-1\right)+Q\end{bmatrix},\end{split}$
@@
In [Chi24], we transform (2.2) using the $XYW$-coordinates, where
@@
Many computations in [Chi24] carry over via (2.11). In particular, the original coordinates are recovered with the same formula. Local solutions that satisfy the initial condition (2.5) is transformed integral curves that emanate $p_{0}:=(1,0,0,0,0,0,0,0,0)$. The positiviteness and finiteness of $b_{0}$ and $c_{0}$ requires the limit
@@
The problem of finding complete Ricci solitons are transformed to finding integral curves emanate $p_{0}$ that are defined on $\mathbb{R}$; see [Chi24, Proposition 2.2].
@@
    $\mathcal{RS}_{m=0}:=\mathcal{RS}_{\text{steady}}\cap\{X_{3},Z_{3}=0\}$
    By (2.12), integral curves on $\mathcal{RS}_{m=0}$ are geometrically degenerate as $(4m+4)$-dimensional Ricci solitons. However, we claim that these integral curves can represent $4$-dimensional steady Ricci solitons on complex line bundle over $\mathbb{CP}^{1}$. To justify this, observe that (2.2) with $m=\epsilon=0$ and $b\equiv c$ is precisely the equation for $4$-dimensional steady Ricci solitons in [App17, (2.1)-(2.3)]. Using the coordinate transformation (2.6), the resulting dynamical system is equivalent to (2.8) on $\mathcal{RS}_{\text{steady}}\cap\mathcal{RS}_{\text{FS}}$ with $m=0$. But such a dynamical system is equivalent to (2.8) on $\mathcal{RS}_{m=0}$. Thus, our claim is justified.
@@ 3. Local Analysis
We apply the $XZW$-coordinates instead of the $XYW$-coordinates in [Chi24] because the linearization at $p_{0}$ has equal positive eigenvalues. This property allows us to apply [CL55, Chapter 13, Theorem 4.5] to analyze functions along integral curves near $p_{0}$.
@@ 4.1. $k\leq 2$
By [Chi24], we have the following lemma.
@@
By [Chi24, Proposition 3.2], the factor $K$ is non-negative on $\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{X_{3}\geq 0\}\cap\{Z_{2}-Z_{3}\geq 0\}$. Finally, by Proposition 4.1, we have $\left.\langle\nabla F_{2},V\rangle\right|_{\mathcal{F}\cap\{F_{2}=0\}}\geq 0$. Therefore, the set $\mathcal{F}$ is compact and invariant. ∎
@@
At $p_{0}$, the inequalities $01-Z_{1}>$ 0 and $X_{1}>0$ hold. The other defining inequalities in (4.5) vanish at $p_{0}$. By (3.12) and (3.16), the defining inequalities
@@ 4.2. $k\geq 3$
By the inequalities $1-Z_{1}\geq 0$ and $Z_{2}-Z_{3}\geq 0$, it is clear that $\mathcal{A}$ and $\mathcal{B}$ are compact. By [Chi24, Proposition 3.2], the set $\mathcal{A}$ is invariant. Since
@@
In other words, if a $\xi(k,\theta,s_{4},s_{5})$ leaves $\mathcal{B}$ at some $\eta_{*}$, either $1-Z_{1}$ or $X_{1}-X_{2}$ vanishes at $\eta_{*}$. These two functions cannot vanish simultaneously as the set $\mathcal{RS}_{\text{round}}$ is invariant while $p_{0}\notin\mathcal{RS}_{\text{round}}$. If the integral curve leaves $\mathcal{B}$ through a point with $01-Z_{1}=$ 0 and $X_{1}-X_{2}>0$, it enters the invariant set $\mathcal{C}$. If the integral curve leaves $\mathcal{B}$ through a point with $01-Z_{1}>$ 0 and $X_{1}-X_{2}=0$, then it enters the invariant set $\mathcal{A}$. The proof is complete. ∎
@@
(4.18) $\begin{split}&\left(Q+1-H+\frac{\epsilon}{2}W(1-H)\right)^{\prime}\\ &=2Q\left(G-\frac{\epsilon}{2}W\right)+\epsilon(H-1)W+2\frac{\epsilon}{2}W\left(G-\frac{\epsilon}{2}W\right)(1-H)\\ &\quad-\left(\frac{\epsilon}{2}W+1\right)\left((H-1)\left(G-\frac{\epsilon}{2}W-1\right)+Q\right)\\ &=\left(Q+1-H+\frac{\epsilon}{2}W(1-H)\right)\left(2G-\frac{3\epsilon}{2}W-1\right)+(1-H)\left(\frac{\epsilon}{2}W-1\right)G.\\ \end{split}$
@@
Initially the integral curve $\xi(k,\theta,s_{4},s_{5})$ is in $\mathcal{B}$. For $s_{4}>0$, the function $Q$ is initially negative. Furthermore, as $R_{i}\geq 0$ in $\mathcal{B}$, we have $-Q\in(0,1)$ along $\xi(k,\theta,s_{4},s_{5})$ in $\mathcal{B}$. We have
@@
(4.28) $\beta_{k,\theta}$ $=\inf\left\{\tilde{\alpha}_{k,\theta}\mid\text{$\xi(k,\theta,s_{4},s_{5})$ does not enter $\mathcal{C}$ for any $s_{4}\geq\tilde{\alpha}_{k,\theta}$ and $s_{5}\geq 0$ }\right\}.$
@@ 5.1. Non-Einstein Ricci solitons
If an integral curve enters $\mathcal{A}$, the analysis in [Chi24, Section 4] applies. We have the following lemma.
@@
We investigate the asymptotic for integral curves that stays in $\mathcal{B}$. The following proposition is extended directly from [Chi24, Section 4].
@@
A $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>0$ and $s_{5}\geq 0$ converges to $(0,0,0,\mu^{2},0,0,0,0)$ for some $\mu\in[0,1]$. The function $Q$ thus converges to $-1$.
@@
The proof of [Chi24, Proposition 4.1] applies verbatim to our case. ∎
@@
If a $\xi(k,\theta,s_{4},0)$ stays in $\mathcal{B}$, the function $Z_{1}$ monotonically increases. Therefore, the limit $\lim\limits_{\eta\to\infty}\sqrt{Z_{1}}=\mu^{2}$ exists for some $\mu\in(0,1]$ by Proposition 5.2. Apply [Win23, Proposition 2.22], the limit $\lim\limits_{\eta\to\infty}\sqrt{\frac{Z_{3}}{Z_{2}}}=\nu^{2}$ also exists. By [Chi24, Lemma 4.7], the possible values for $(\mu^{2},\nu^{2})$ are $(1,1)$ and $\left(1,\frac{1}{2m+3}\right)$, and we have the limits
@@
as $t\to\infty$. The paraboloid is the standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=1$) or the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=\frac{1}{2m+3}$). If $\theta=0$, the integral curve $\xi(k,0,s_{4},0)$ stays in $\mathcal{RS}_{\text{FS}}$, where $X_{2}=X_{3}$ and $Z_{2}=Z_{3}$. The base of the asymptotic paraboloid is round. Since $\lim\limits_{\eta\to-\infty}\frac{Z_{3}}{Z_{2}}<1$ initially along $\xi(k,\theta,s_{4},0)$ with $\theta>0$, the limit $\nu^{2}\neq 1$ by [Chi24, Proposition 4.9]. The base of the asymptotic paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. ∎

### v2

@@ Abstract.
We construct a continuous 3-parameter family of non-shrinking Ricci solitons on complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{2m+1}$, where the base space is not necessarily Kähler–Einstein. Each $\mathcal{O}(k)$ with $k\in[3,2m+1]$ admits at least one asymptotically conical (AC) Ricci-flat metric in this family. For each $\mathcal{O}(k)$ with $k\geq 3$, the family includes infinitely many asymptotically paraboloidal (AP) steady Ricci solitons.
2020 Mathematics Subject Classification: 53E20 (primary), 53C25 (secondary). Keywords: Ricci soliton, cohomogeneity one metric. The author is supported by the NSFC (No. 12301078), the Natural Science Foundation of Jiangsu Province (BK-20220282), and XJTLU Research Development Fund (RDF-21-02-083).
@@ 1. Introduction
Ricci solitons are candidates for blow-up singularities of the Ricci flow. Specifically, a shrinking Ricci soliton is a candidate for a Type I singularity while a steady Ricci soliton is a candidate for a Type II singularity. From Perelman’s no-collapsing theorem [Per02], Type II singularities must be non-collapsed.
Complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{n}$ provide a rich source of examples for gradient Ricci solitons. The principal orbit of $\mathcal{O}(k)$ is the lens space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$, a circle bundle over $\mathbb{CP}^{n}$. The integer $k$ classifies the circle bundle, corresponding to a multiple of the indivisible element in $H^{2}(\mathbb{CP}^{n},\mathbb{Z})$. This setup has led to numerous examples of Ricci solitons with various asymptotic behaviors. The Bérard-Bergery metric on $\mathcal{O}(n+1)$ is an AC Ricci-flat metric; see [BB82]. By [Cao96], it is known that $\mathcal{O}(n+1)$ admits a 1-parameter family of steady Kähler–Ricci solitons. These solitons have collapsed volume growth. In particular, they are asymptotically cigar-paraboloidal (ACP). Expanding Kähler–Ricci solitons on $\mathcal{O}(k)$ with $k>n+1$ are found in [FIK03]. The $U(n+1)$-invariant ansatz is further generalized to complex line bundles over products of Fano manifolds in [DW11]. Recently, new steady Ricci solitons were found on $\mathcal{O}(k)$ independently by [Win23], [Sto24] and [App17]. It was proved in [App17] that each $\mathcal{O}(k)$ with $k>n+1$ admits a non-collapsed AP steady Ricci soliton.
@@
The group action descends to $\mathbb{CP}^{2m+1}$, where the isotropy subgroup of [0::0:1][0:...:0:1] is
@@
Since the principal orbit is locally isometric to $\mathbb{S}^{4m+3}$, the cohomogeneity one dynamical system on $\mathcal{O}(k)$ is the same for different values of $k$. The same dynamical system also appeared in the author’s earlier work [Chi26], where a continuous 1-parameter family of non-collapsed steady Ricci solitons were constructed on Euclidean spaces. However, the geometric interpretation of the solutions is different in the present setting. Here we study complex line bundles $\mathcal{O}(k)\to\mathbb{CP}^{2m+1}$, where the global existence of complete metrics depends crucially on the integer parameter $k$. This leads to new families of complete Ricci solitons.
For $k\in\{1,2\}$, we construct a compact invariant set to obtain the following theorem.
@@
On each $\mathcal{O}(k)$ with $3\leq k\leq 2m+1$, there exists at least one AC Ricci-flat metric, whose base of the cone is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. On each $\mathcal{O}(k)$ with $k\geq 3$, there exist infinitely many AP steady Ricci solitons whose base of the paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$.
This article is structured as the following. We begin by presenting the cohomogeneity one Ricci soliton equation and introducing a new coordinate transformation, differing from that in [Chi26], to establish local existence. For global analysis, we first construct an invariant compact set to prove Theorem 1.1. Next, we generalize [App17, Theorem 5.1] and prove Theorem 1.2-1.3. Finally, we investigate the asymptotics for these Ricci solitons and prove Theorem 1.4.
@@ 2. Cohomogeneity one equation
The geometric setup is similar to the one in [Chi26]. Specifically, let $Q$ be the inner product on $\mathfrak{g}$ that generates the round metric on $\mathbb{S}^{4m+3}$. We have a $Q$-orthogonal decomposition of the isotropy representation $\mathfrak{g}/\mathfrak{k}$ into three irreducible summands. Consider the ansatz
@@
We set $f(0)=0$ without loss of generality. By [Buz11], the potential function $f$ is an even function in $t$ near 0. As the $\mathbb{S}^{1}$-fiber collapses at the singular orbit, we have $\mathrm{tr}(L)=\frac{1}{t}+O(t)$ as $t\to 0$. Therefore, we have $2\ddot{f}(0)=C$ by(2.3). As the radius for the $\mathbb{S}^{1}$-fiber of $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ is $\frac{1}{k}$, we require $\dot{a}(0)=k$ for smooth extension and we consider the initial condition
@@
[…]
@@
In [Chi26], we transform (2.2) using the $XYW$-coordinates, where
@@
Many computations in [Chi26] carry over via (2.11). In particular, the original coordinates are recovered with the same formula. Local solutions that satisfy the initial condition (2.5) is transformed integral curves that emanate $p_{0}:=(1,0,0,0,0,0,0,0,0)$. The positiviteness and finiteness of $b_{0}$ and $c_{0}$ requires the limit
@@
The problem of finding complete Ricci solitons are transformed to finding integral curves emanate $p_{0}$ that are defined on $\mathbb{R}$; see [Chi26, Proposition 2.2].
@@
    $\mathcal{RS}_{m=0}:=\mathcal{RS}_{\text{steady}}\cap\{X_{3},Z_{3}=0\}$.
    Although integral curves in $\mathcal{RS}_{m=0}$ are geometrically degenerate when viewed as $(4m+4)$-dimensional Ricci solitons, they can be interpreted as four-dimensional steady Ricci solitons on complex line bundles over $\mathbb{CP}^{1}$. Indeed, setting $m=\epsilon=0$ in (2.2) and ignoring the function $c$ yields precisely the system for four-dimensional steady Ricci solitons as in [App17, (2.1)–(2.3)]. Applying the coordinate transformation (2.6), this system is equivalent to the restriction of (2.8) to $\mathcal{RS}_{\text{steady}}\cap\mathcal{RS}_{\text{FS}}$ with $m=0$, which in turn coincides with the dynamical system on $\mathcal{RS}_{m=0}$. It is important to note that, unlike the four-dimensional expanders studied in [Wan25, Don25], this $m=0$ subsystem produces only steady solitons.
@@ 3. Local Analysis
We apply the XZWXZW-coordinates instead of the XYWXYW-coordinates in [Chi26] because the linearization at $p_{0}$ has equal positive eigenvalues. This property allows us to apply [CL55, Chapter 13, Theorem 4.5] to analyze functions along integral curves near $p_{0}$.
@@ 4.1. $k\leq 2$
By [Chi26], we have the following lemma.
@@
By [Chi26, Proposition 3.2], the factor $K$ is non-negative on $\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{X_{3}\geq 0\}\cap\{Z_{2}-Z_{3}\geq 0\}$. Finally, by Proposition 4.1, we have $\left.\langle\nabla F_{2},V\rangle\right|_{\mathcal{F}\cap\{F_{2}=0\}}\geq 0$. Therefore, the set $\mathcal{F}$ is compact and invariant. ∎
@@
At $p_{0}$, the inequalities $1-Z_{1}>0$ and $X_{1}>0$ hold. The other defining inequalities in (4.5) vanish at $p_{0}$. By (3.12) and (3.16), the defining inequalities
@@ 4.2. $k\geq 3$
By the inequalities $1-Z_{1}\geq 0$ and $Z_{2}-Z_{3}\geq 0$, it is clear that $\mathcal{A}$ and $\mathcal{B}$ are compact. By [Chi26, Proposition 3.2], the set $\mathcal{A}$ is invariant. Since
@@
In other words, if a $\xi(k,\theta,s_{4},s_{5})$ leaves $\mathcal{B}$ at some $\eta_{*}$, either $1-Z_{1}$ or $X_{1}-X_{2}$ vanishes at $\eta_{*}$. These two functions cannot vanish simultaneously as the set $\mathcal{RS}_{\text{round}}$ is invariant while $p_{0}\notin\mathcal{RS}_{\text{round}}$. If the integral curve leaves $\mathcal{B}$ through a point with $1-Z_{1}=0$ and $X_{1}-X_{2}>0$, it enters the invariant set $\mathcal{C}$. If the integral curve leaves $\mathcal{B}$ through a point with $1-Z_{1}>0$ and $X_{1}-X_{2}=0$, then it enters the invariant set $\mathcal{A}$. The proof is complete. ∎
@@
[…]
@@
Initially the integral curve $\xi(k,\theta,s_{4},s_{5})$ is in $\mathcal{B}$. For $s_{4}>0$, the function $Q$ is initially negative. Furthermore, as $R_{i}\geq 0$ in $\mathcal{B}$, we have $Q\in(0,1)$ along $\xi(k,\theta,s_{4},s_{5})$ in $\mathcal{B}$. We have
@@
[…]
@@ 5.1. Non-Einstein Ricci solitons
If an integral curve enters $\mathcal{A}$, the analysis in [Chi26, Section 4] applies. We have the following lemma.
@@
We investigate the asymptotic for integral curves that stays in $\mathcal{B}$. The following proposition is extended directly from [Chi26, Section 4].
@@
A $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>0$ and $s_{5}\geq 0$ converges to $(0,0,0,\mu^{2},0,0,0,0)$ for some $\mu\in[0,1]$. The function $Q$ thus converges to 1-1.
@@
The proof of [Chi26, Proposition 4.1] applies verbatim to our case. ∎
@@
If a $\xi(k,\theta,s_{4},0)$ stays in $\mathcal{B}$, the function $Z_{1}$ monotonically increases. Therefore, the limit $\lim\limits_{\eta\to\infty}\sqrt{Z_{1}}=\mu^{2}$ exists for some $\mu\in(0,1]$ by Proposition 5.2. Apply [Win23, Proposition 2.22], the limit $\lim\limits_{\eta\to\infty}\sqrt{\frac{Z_{3}}{Z_{2}}}=\nu^{2}$ also exists. By [Chi26, Lemma 4.7], the possible values for $(\mu^{2},\nu^{2})$ are $(1,1)$ and $\left(1,\frac{1}{2m+3}\right)$, and we have the limits
@@
as $t\to\infty$. The paraboloid is the standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=1$) or the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=\frac{1}{2m+3}$). If $\theta=0$, the integral curve $\xi(k,0,s_{4},0)$ stays in $\mathcal{RS}_{\text{FS}}$, where $X_{2}=X_{3}$ and $Z_{2}=Z_{3}$. The base of the asymptotic paraboloid is round. Since $\lim\limits_{\eta\to-\infty}\frac{Z_{3}}{Z_{2}}<1$ initially along $\xi(k,\theta,s_{4},0)$ with $\theta>0$, the limit $\nu^{2}\neq 1$ by [Chi26, Proposition 4.9]. The base of the asymptotic paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. ∎